OpenStudy (anonymous):

Mary and her brother John collect foreign coins. Mary has three times the number of coins that John has. Together they have 140 foreign coins. Find how many coins Mary has.

OpenStudy (ganpat):

M = 3J M + J = 140.. substitute value of M 3J + J = 140 4J = 140 J = 35.. M = 3 * 35 = 105 foreign currencies...

OpenStudy (anonymous):

Thanks a lot

OpenStudy (ganpat):

w/c..